An aluminum block with mass m was nudged, it slides down the inclined surface at a constant speed when the angle of inclination is 220. Determine the coefficient of kinetic friction between the aluminum block and the inclined surface.


1
Expert's answer
2020-07-20T14:54:03-0400


Here "\mathbf{mg}" is the gravity force, "\mathbf{F_{fr}}" is the friction force and "\mathbf{N}" is the reactance,

Since the block is moving with the constant speed, its acceleration is zero. Thus, writting down Newton's second law in the projections on axes, obtain

x-axis:


"mg\sin\theta - F_{fr} = 0 \\
F_{fr} = mg\sin\theta"

On the other hand,


"F_{fr} = \mu N"

where "\mu" is the coefficient of kinetic friction between the aluminum block and the inclined surface.

Thus, get:


"\mu N = mg\sin\theta\\
\mu = \dfrac{mg\sin\theta}{N}"

y-axis:


"N-mg\cos\theta = 0\\
N = mg\cos\theta"

Substituting "N" into the "\mu", obtain:


"\mu = \dfrac{mg\sin\theta}{mg\cos\theta } = \tan\theta\\
\mu = \tan22\degree \approx 0.404"

Answer. 0.404.